Hale Corporation is comparing two different capital structures, an all-equity plan (Plan I) and a levered plan (Plan II). Under Plan I, the company would have 180,000 shares of stock outstanding. Under Plan II, there would be 130,000 shares of stock outstanding and $2.27 million in debt outstanding. The interest rate on the debt is 8 percent and there are no taxes.

Use MM Proposition I to find the price per share.

What is the value of the firm under each of the two proposed plans?

a) all equity plan

b) levered plan
Answer #1

a) From the description (180,000-130,000) shares = 2.27 million
Value of 50,000 shares = 2..27 million
Value of 180,000 shares = 2.27*180,000/50,000 = 8.172 million
Value of firm = 8.172 million
Price per share = 8,172,000/180,000 = 45.4

b) Value of Levered firm = Value of unlevered form if there are no taxes = 8.172
Price per share of Levered firm = (8,172,000-2,270,000)/130,000 = 45.4

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well